Andy is 5 years older than Chris in 3 years the sum of their ages will be 35. what are their ages
Answer by aaronwiz(69)

Hi, my name is Aaron I am in 10th grade and im in honors trig/pre-calc. what you are going to do is make a system of equations..

a=andy
c=chris
c+5=a
(c+3)+(a+3)=35
simplify it down. so

a-c=5
a+c=29
now just solve, which im sure you can do